# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

### v2

@@ 5.3 Variation of parameter in the wind shells
Up to this point, the discussion has regarded observational trends along the direction of the wind streamlines, indicating an overall agreement with a DW scenario.
@@
However, the most puzzling evidence that emerges from the data is the inhomogeneity of the wind across its width, as indicated by the presence of distinct nested shells with enhanced emission. This property, not expected in a standard self-similar DW, is actually found in many molecular outflows observed at the high spatial and spectral resolution offered by ALMA (e.g. Zhang et al. 2019; Fernández-López et al. 2020; López-Vázquez et al. 2024; Ai et al. 2024; de Valon et al. 2020, 2022; Omura et al. 2024; Nazari et al. 2024, cf. Introduction).
@@
The analysis of the HL Tau wind offers the advantage of allowing a comparison between the structures in the wind and in the disk. The derived configuration of the wind shells, the large estimated footpoint radii, and the indication of origin of the shells from denser gas rings of the disk appear to support recent models that prescribe the spontaneous formation of a ring-gap substructure in the disk and of a connected inhomogeneous wind, as a result of magnetic instabilities (Béthune et al., 2017; Suriano et al., 2018, 2019; Riols & Lesur, 2019; Riols et al., 2020; Cui & Bai, 2021; Hu et al., 2022). In these numerical studies the secular evolution of protoplanetary disks is studied including ambipolar diffusion in the generalized Ohm’s equation. As a result, one sees the rapid development of a self-organization of the magnetic field in the disk, which accumulates in the regions of lower mass density. In this way a system of rings and gaps forms quickly in the disk, with a greater concentration of magnetic field in the gaps. The final configuration also sees the formation of a transversely inhomogeneous disk wind arranged in nested shells, with alternate layers of enhanced density originating from the gaseous disk rings and layers of faster lighter flow originating from the highly magnetized gaps. The ALMA observations like the ones described in this paper provide maps of the intensity of the line emission, and hence are prone to detect the wind layers of higher density, that is, in the above scenario, the layers arising from the disk rings.
@@
The interest of this approach is the alternative it offers to models in which the disk substructure is due to the presence of young planets. Given the lack of detection of planetary bodies in the HL Tau disk (Mullin et al., 2024) as well as in the other structured disks (with the exception of PDS 70, where, however, the planets reside in a very large cavity, Keppler et al. 2018), the point is of crucial importance, as it indicates that magnetic effects must be taken into account in the investigations aimed at the determination of the timescale of formation of the planets themselves.
@@
However, a number of issues need to be addressed before we can confirm the validity of the proposed association. First of all, the value of the magnetic lever arm assumed in this class of numerical models is generally 2. A similar simulation for , as suggested by our analysis, would be desirable. Secondly, the vertical extension of the structuration of the flow varies among the models, but in general the simulations do not reach the region at 103 au from the star in which we observe the wind structures with ALMA. An extension of numerical studies to large vertical scales would be needed, but it would come at the expense of a consistent computational effort. Finally, in these simulations, the disk structures are often seen to slowly move (Suriano et al., 2019; Martel & Lesur, 2022), questioning the survival of their connection to the wind. However, the motion occurs on secular timescales, suggesting that the coherence of the magnetic flux tubes is probably not disrupted. Clearly, the applicability of the class of models described above to the observed cases needs to be tested with refined theoretical analyses, which, however, are out of the scope of this work.
@@ 6 Conclusions
We investigated the properties of the molecular CO wind associated with the young HL Tau system by means of a detailed spectro-imaging analysis of the 12CO (2-1) line emission, observed at 03 angular resolution in the context of the ALMA-DOT campaign (Podio et al., 2020a; Garufi et al., 2021). Our analysis confirms previous results derived from integrated intensity maps, but new insights on the structure of the flow come from the inspection of the individual projections of the Position-Position-Velocity (PPV) ALMA datacube in the brighter and less perturbed red-shifted lobe.
@@
A complex substructure is revealed in the velocity channel maps, constituted by concatenated bubble- and arc-shaped features that progressively increase in size and distance from the source with velocity. The conical shape of the wind appears to be the result of superposition of the limb-brightened borders of these features. The position-velocity (PV) diagram formed along the system axis presents a fan of separated almost linear traces opening from the source region, and with increasing velocity with offset from the star. The traces have strict spatial correspondence with the arcs in the channel maps. Finally, PV diagrams taken across the flow at stepped separations from the source show a series of concatenated ellipse-shaped features that correspond directly to the traces in the longitudinal PV diagram and in the channel maps. The ellipses are tilted with respect to the system axis, indicating rotation of the flow in clockwise direction looking from the SW lobe toward the star, which matches the disk rotation sense.
@@
On the basis of these projections we infer a global structure of the emission in the PPV datacube consisting of a series of distinct concatenated paraboloid-like surfaces with the apex toward the source, with progressively larger aperture as the angle of their axis with respect to the velocity axis increases (cf. Fig. 7). In the 3D physical space this translates into a finite number of coaxial, distinct, rotating nested gas shells. Ellipse fitting of the three directly identifiable structures in adjacent transverse PV diagrams allows a tomographic reconstruction of the morphology and kinematics of the outer three shells of the flow. We find smaller radii, higher poloidal and toroidal velocities, and steeper increases of the velocity with distance from the source for shells progressively closer to the system axis.
@@
This structure can be reproduced by a number of different wind acceleration mechanisms. In this paper, we focus on an analysis of the applicability of magnetohydrodynamic (MHD) disk winds (DW) accelerated from an extended region of the disk surface. We defer the comparison with other classes of models to future investigations.
@@
Under the DW hypothesis, we derive from the rotation properties of the flow that each examined shell is rooted in the disk within a few au of one of three adjacent disk rings, seen in the continuum emission of the dust, located at 58, 72 and 86 au from the star (ALMA Partnership et al., 2015). However, we caution that, because of the faintness of the examined flow substructures, the uncertainty on the footpoint radii is greater than the distance of the adjacent rings and gaps.
@@
We attempted a determination of the magnetic lever arm in the wind by comparing the observed trends of the wind parameters along the shells with the predictions of models of magnetized winds from either a laminar AD-dominated disk or a fully turbulent disk. The characteristics of the HL Tau wind turn out to be intermediate between the two types of solutions, with a value of . We discuss this finding in the context of the characteristics of the examined models. The retrieved launching region between 50 and 90 au would point to a wind from an AD-dominated outer disk, but solutions with such a high value of do not exist and have yet to be found (current solutions reach = 2.76). A possibility would be to consider solutions for a higher level of AD than is usually assumed, which may be applicable if the wind itself screens the incoming ionizing cosmic rays by magnetic mirroring (Cleeves et al., 2015). An increased field amplitude in the outer disk region with respect to the one commonly adopted could also help justify the high value of , as the ejection efficiency in the wind would be higher.
@@
A magnetically driven disk wind with the above characteristics could sustain disk accretion at large scale by removing angular momentum in regions where the effective turbulence is suppressed because of weak internal ionization. In addition, the presence of distinct denser nested wind shells rooted in disk rings appears to support the results of recent non-ideal MHD numerical simulations for AD-dominated disks, which see the spontaneous formation of a ring-gap substructure in the disk connected with an inhomogeneous layered flow (e.g. Suriano et al. 2019; Riols et al. 2020), in a way alternative to the action of the yet elusive protoplanets. These aspects may have profound implications for the evolution of the disk and the process of planet formation.
@@
The inferred properties of the CO wind call for further analyses of both observational and theoretical type. An analogous scrutiny of the fainter components and substructures in both the red- and blue-shifted lobes in the same datacube, as well as the challenging estimate of the mass and angular momentum transported in each shell, requires the development of dedicated procedures, which is currently underway (Nony et al., in prep.). On the theoretical side, it would be interesting to consider the extension of the numerical simulations of inhomogeneous layered disk winds to the large scales probed by our observations. Then, a comparison of the retrieved results with other classes of wind models is to be attempted, and it will the subject of future works. Finally, new observational studies for this and other systems, currently in progress, will allow us to analyze the dynamic relationship between the outer CO outflow, the coaxial H2 wind and the atomic jet.
@@
These investigations are expected to provide further insight on the nature of the ejection phenomenon and finally to make us able to quantify the feedback of the winds on the disk evolution.
@@ Acknowledgements.
This paper uses ALMA data 2018.1.01037.S (PI L. Podio). ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. We thank the anonymous referee for her/his precious comments that led to improved results. The authors are grateful to G. Lesur for fruitful discussions and for his support in using the self-similar MHD disk wind solutions that he has made available to the public at: https://github.com/glesur/PPDwind. The authors thank Z.-Y. Li, Y. Bai, T. Downes, F. Louvet, D. Fedele, and L. Testi for their helpful comments, and K. Rygl, V. Delabrosse, and U. Locatelli for their help with data processing and analysis. This work was supported by Large Grant INAF 2022 ’YSOs Outflows, Disks, and Accretion: towards a global framework for the evolution of planet-forming systems (YODA)’. Part of the research activities were carried out with the contribution of the Next Generation EU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 - From Research to Business (M4C2), Investment Line 3.1 - Strengthening and creation of Research Infrastructures, Project IR0000034 – ”STILES - Strengthening the Italian Leadership in ELT and SKA”. LP and CC acknowledge the projects PRIN-MUR 2020 - ”BEYOND-2p: Astrochemistry beyond the second period elements” (Prot. 2020AFB3FX); ASI-Astrobiologia 2023 - ”MIGLIORA-Modeling Chemical Complexity”, (F83C23000800005); INAF-GO 2023 - ”PROTO-SKA-Exploiting ALMA data to study planet-forming disks: preparing the advent of SKA” (C13C23000770005); INAF MiniGrant 2022 - ”Chemical Origins” (PI: L. Podio); National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 - Next Generation EU – 2022JC2Y93 ”Chemical Origins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks” (CUP J53D23001600006) - Grant Assignment Decree No. 962 adopted on 30.06.2023 by the Italian Ministry of University and Research.References
- Ai et al. (2024) Ai, T.-H., Liu, C.-F., Shang, H., Johnstone, D., & Krasnopolsky, R. 2024, ApJ, 964, 147
- Alexander et al. (2014) Alexander, R., Pascucci, I., Andrews, S., Armitage, P., & Cieza, L. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 475–496
- ALMA Partnership et al. (2015) ALMA Partnership, Brogan, C. L., Pérez, L. M., et al. 2015, ApJ, 808, L3
- Anderson et al. (2003) Anderson, J. M., Li, Z.-Y., Krasnopolsky, R., & Blandford, R. D. 2003, ApJ, 590, L107
- Andrews et al. (2018) Andrews, S. M., Huang, J., Pérez, L. M., et al. 2018, ApJ, 869, L41
- Anglada et al. (2007) Anglada, G., López, R., Estalella, R., et al. 2007, AJ, 133, 2799
- Bacciotti et al. (2002) Bacciotti, F., Ray, T. P., Mundt, R., Eislöffel, J., & Solf, J. 2002, ApJ, 576, 222
- Beck et al. (2008) Beck, T. L., McGregor, P. J., Takami, M., & Pyo, T.-S. 2008, ApJ, 676, 472
- Béthune et al. (2017) Béthune, W., Lesur, G., & Ferreira, J. 2017, A&A, 600, A75
- Bjerkeli et al. (2016) Bjerkeli, P., van der Wiel, M. H. D., Harsono, D., Ramsey, J. P., & Jørgensen, J. K. 2016, Nature, 540, 406
- Blandford & Payne (1982) Blandford, R. D. & Payne, D. G. 1982, MNRAS, 199, 883
- Casse & Ferreira (2000a) Casse, F. & Ferreira, J. 2000a, A&A, 353, 1115
- Casse & Ferreira (2000b) Casse, F. & Ferreira, J. 2000b, A&A, 361, 1178
- Cleeves et al. (2015) Cleeves, L. I., Bergin, E. A., Qi, C., Adams, F. C., & Öberg, K. I. 2015, ApJ, 799, 204
- Codella et al. (2020) Codella, C., Podio, L., Garufi, A., et al. 2020, A&A, 644, A120
- Cui & Bai (2021) Cui, C. & Bai, X.-N. 2021, MNRAS, 507, 1106
- De Valon (2021) De Valon, A. 2021, PhD thesis, Universite de Grenoble Alpes
- de Valon et al. (2020) de Valon, A., Dougados, C., Cabrit, S., et al. 2020, A&A, 634, L12
- de Valon et al. (2022) de Valon, A., Dougados, C., Cabrit, S., et al. 2022, A&A, 668, A78
- Delabrosse et al. (2024) Delabrosse, V., Dougados, C., Cabrit, S., et al. 2024, A&A, 688, A173
- Fernández-López et al. (2020) Fernández-López, M., Zapata, L. A., Rodríguez, L. F., et al. 2020, AJ, 159, 171
- Ferreira (1997) Ferreira, J. 1997, A&A, 319, 340
- Ferreira et al. (2006) Ferreira, J., Dougados, C., & Cabrit, S. 2006, A&A, 453, 785
- Flock et al. (2019) Flock, M., Turner, N. J., Mulders, G. D., et al. 2019, A&A, 630, A147
- Frank et al. (2014) Frank, A., Ray, T. P., Cabrit, S., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 451–474
- Galli et al. (2018) Galli, P. A. B., Loinard, L., Ortiz-Léon, G. N., et al. 2018, ApJ, 859, 33
- Garufi et al. (2021) Garufi, A., Podio, L., Codella, C., et al. 2021, A&A, 645, A145
- Garufi et al. (2020) Garufi, A., Podio, L., Codella, C., et al. 2020, A&A, 636, A65
- Garufi et al. (2022) Garufi, A., Podio, L., Codella, C., et al. 2022, A&A, 658, A104
- Hu et al. (2022) Hu, X., Li, Z.-Y., Zhu, Z., & Yang, C.-C. 2022, MNRAS, 516, 2006
- Jacquemin-Ide et al. (2019) Jacquemin-Ide, J., Ferreira, J., & Lesur, G. 2019, MNRAS, 490, 3112
- Keppler et al. (2018) Keppler, M., Benisty, M., Müller, A., et al. 2018, A&A, 617, A44
- Klaassen et al. (2016) Klaassen, P. D., Mottram, J. C., Maud, L. T., & Juhasz, A. 2016, MNRAS, 460, 627
- Lee et al. (2017) Lee, C.-F., Ho, P. T. P., Li, Z.-Y., et al. 2017, Nature Astronomy, 1, 0152
- Lee et al. (2018) Lee, C.-F., Li, Z.-Y., Codella, C., et al. 2018, ApJ, 856, 14
- Lee et al. (2000) Lee, C.-F., Mundy, L. G., Reipurth, B., Ostriker, E. C., & Stone, J. M. 2000, ApJ, 542, 925
- Lesur et al. (2023) Lesur, G., Flock, M., Ercolano, B., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 465
- Lesur (2021) Lesur, G. R. J. 2021, A&A, 650, A35
- Lodato et al. (2019) Lodato, G., Dipierro, G., Ragusa, E., et al. 2019, MNRAS, 486, 453
- López-Vázquez et al. (2024) López-Vázquez, J. A., Lee, C.-F., Fernández-López, M., et al. 2024, ApJ, 962, 28
- Louvet et al. (2018) Louvet, F., Dougados, C., Cabrit, S., et al. 2018, A&A, 618, A120
- Lumbreras & Zapata (2014) Lumbreras, A. M. & Zapata, L. A. 2014, AJ, 147, 72
- Martel & Lesur (2022) Martel, É. & Lesur, G. 2022, A&A, 667, A17
- McMullin et al. (2007) McMullin, J. P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R. A. Shaw, F. Hill, & D. J. Bell, 127
- Movsessian et al. (2012) Movsessian, T. A., Magakian, T. Y., & Moiseev, A. V. 2012, A&A, 541, A16
- Müller et al. (2001) Müller, H. S. P., Thorwirth, S., Roth, D. A., & Winnewisser, G. 2001, A&A, 370, L49
- Mullin et al. (2024) Mullin, C., Dong, R., Leisenring, J., et al. 2024, AJ, 167, 183
- Mundt et al. (1990) Mundt, R., Buehrke, T., Solf, J., Ray, T. P., & Raga, A. C. 1990, A&A, 232, 37
- Nazari et al. (2024) Nazari, P., Tabone, B., Ahmadi, A., et al. 2024, A&A, 686, A201
- Nisini et al. (2024) Nisini, B., Navarro, M. G., Giannini, T., et al. 2024, ApJ, 967, 168
- Omura et al. (2024) Omura, M., Tokuda, K., & Machida, M. N. 2024, ApJ, 963, 72
- Pascucci et al. (2025) Pascucci, I., Beck, T. L., Cabrit, S., et al. 2025, Nature Astronomy, 9, 81
- Pascucci et al. (2023) Pascucci, I., Cabrit, S., Edwards, S., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 567
- Pinte et al. (2016) Pinte, C., Dent, W. R. F., Ménard, F., et al. 2016, ApJ, 816, 25
- Pinto et al. (2008) Pinto, C., Galli, D., & Bacciotti, F. 2008, A&A, 484, 1
- Podio et al. (2019) Podio, L., Bacciotti, F., Fedele, D., et al. 2019, A&A, 623, L6
- Podio et al. (2011) Podio, L., Eislöffel, J., Melnikov, S., Hodapp, K. W., & Bacciotti, F. 2011, A&A, 527, A13
- Podio et al. (2020a) Podio, L., Garufi, A., Codella, C., et al. 2020a, A&A, 642, L7
- Podio et al. (2020b) Podio, L., Garufi, A., Codella, C., et al. 2020b, A&A, 644, A119
- Podio et al. (2021) Podio, L., Tabone, B., Codella, C., et al. 2021, A&A, 648, A45
- Pyo et al. (2006) Pyo, T.-S., Hayashi, M., Kobayashi, N., et al. 2006, ApJ, 649, 836
- Rabenanahary et al. (2022) Rabenanahary, M., Cabrit, S., Meliani, Z., & Pineau des Forêts, G. 2022, A&A, 664, A118
- Ray & Ferreira (2021) Ray, T. P. & Ferreira, J. 2021, New A Rev., 93, 101615
- Riols & Lesur (2019) Riols, A. & Lesur, G. 2019, A&A, 625, A108
- Riols et al. (2020) Riols, A., Lesur, G., & Menard, F. 2020, A&A, 639, A95
- Segura-Cox et al. (2020) Segura-Cox, D. M., Schmiedeke, A., Pineda, J. E., et al. 2020, Nature, 586, 228
- Shang et al. (2020) Shang, H., Krasnopolsky, R., Liu, C.-F., & Wang, L.-Y. 2020, ApJ, 905, 116
- Shang et al. (2023) Shang, H., Liu, C.-F., Krasnopolsky, R., & Wang, L.-Y. 2023, ApJ, 944, 230
- Sheehan & Eisner (2017) Sheehan, P. D. & Eisner, J. A. 2017, ApJ, 840, L12
- Sheehan & Eisner (2018) Sheehan, P. D. & Eisner, J. A. 2018, ApJ, 857, 18
- Stephens et al. (2023) Stephens, I. W., Lin, Z.-Y. D., Fernández-López, M., et al. 2023, Nature, 623, 705
- Suriano et al. (2018) Suriano, S. S., Li, Z.-Y., Krasnopolsky, R., & Shang, H. 2018, MNRAS, 477, 1239
- Suriano et al. (2019) Suriano, S. S., Li, Z.-Y., Krasnopolsky, R., Suzuki, T. K., & Shang, H. 2019, MNRAS, 484, 107
- Tabone et al. (2017) Tabone, B., Cabrit, S., Bianchi, E., et al. 2017, A&A, 607, L6
- Tabone et al. (2020) Tabone, B., Cabrit, S., Pineau des Forêts, G., et al. 2020, A&A, 640, A82
- Tabone et al. (2018) Tabone, B., Raga, A., Cabrit, S., & Pineau des Forêts, G. 2018, A&A, 614, A119
- Takami et al. (2007) Takami, M., Beck, T. L., Pyo, T.-S., McGregor, P., & Davis, C. 2007, ApJ, 670, L33
- Welch et al. (2000) Welch, W. J., Hartmann, L., Helfer, T., & Briceño, C. 2000, ApJ, 540, 362
- Yen et al. (2019a) Yen, H.-W., Gu, P.-G., Hirano, N., et al. 2019a, ApJ, 880, 69
- Yen et al. (2019b) Yen, H.-W., Takakuwa, S., Gu, P.-G., et al. 2019b, A&A, 623, A96
- Zapata et al. (2015) Zapata, L. A., Lizano, S., Rodríguez, L. F., et al. 2015, ApJ, 798, 131
- Zhang et al. (2016) Zhang, Y., Arce, H. G., Mardones, D., et al. 2016, ApJ, 832, 158
- Zhang et al. (2019) Zhang, Y., Arce, H. G., Mardones, D., et al. 2019, ApJ, 883, 1
- Zimniak et al. (2024) Zimniak, N., Ferreira, J., & Jacquemin-Ide, J. 2024, A&A, 692, A99
Appendix A Atlas of velocity channel maps
Figures 14 and 15 present an atlas of channel maps of CO (2-1) emission at VLSR from -3.0 km sto +28.6 km s, in steps of 0.4 km s. The panels at VLSR from +3.0 to +8.6 km sare centered on the source, to visualize the low-velocity emission in both lobes, while outside this range, emission is detected only in either the blue- or in the red-shifted lobe, so the maps cover the relevant quadrant, with the source positioned in one corner.
The emission in the blue-shifted NE lobe is very faint. It consists of a large arc-shaped feature opening toward the NE seen at progressively larger distances from the source with increasing blue-shifted velocities. A finger-like feature is identified at almost all velocities in about the same position north to the source and appears to be independent on the moving arc.
The red-shifted SW lobe is distributed in two main components, a low-velocity wide-angle wind and a faster inner flow of apparent conical shape, with an opening angle decreasing with velocity. To quantify the variation of the aperture, we determined, for selected channels from 9 to 28 km sand separately for the SE and NW sides, the angle between the system axis and a line connecting the source and the position of the peak emission at the flow borders (identified by the contour profiles), at an offset of 45 and 75 from the source. The results of this analysis are shown in Figure 16. The half-opening angle decreases on average with velocity, from about 40 2 at 10 km sto 27 2 at 26 km s. The difference in value displayed in a given channel by the various curves reflects the deviations in the shape of the flow from a perfect cone.

### v3

@@ 5.5 Implications of the observed shell separation
Up to this point, the discussion regarded observational trends along the direction of the wind streamlines, indicating an overall agreement with a DW scenario.
@@
However, the most puzzling evidence that emerges from the data is the inhomogeneity of the wind across its width, as indicated by the presence of distinct nested shells with enhanced emission. This property, not expected in a standard self-similar DW, is actually found in many molecular outflows observed at the high spatial and spectral resolution offered by ALMA (e.g., Zhang19; Fernandez20; Lopez24; Ai24; DeValon20; DeValon22; Omura24; Nazari24, cf. Introduction).
@@
The analysis of the HL Tau wind offers the advantage of a comparison between the structures in the wind and in the disk. The derived configuration of the wind shells, the large estimated footpoint radii, and the indication of origin of the shells from denser gas rings of the disk appear to support recent models that prescribe the spontaneous formation of a ring–gap substructure in the disk and of a connected inhomogeneous wind, as a result of magnetic instabilities (Bethune17; Suriano18; Suriano19; Riols19; Riols20; Cui21; Hu22). In these numerical studies the secular evolution of protoplanetary disks is studied including ambipolar diffusion in the generalized Ohm’s equation. As a result, one sees the rapid development of a self-organization of the magnetic field in the disk, which accumulates in the regions of lower mass density. In this way a system of rings and gaps forms quickly in the disk, with a greater concentration of magnetic field in the gaps. The final configuration also sees the formation of a transversely inhomogeneous disk wind arranged in nested shells, with alternate layers of enhanced density originating from the gaseous disk rings and layers of faster lighter flow originating from the highly magnetized gaps. The ALMA observations like the ones described in this paper provide maps of the intensity of the line emission, and hence are prone to detect the wind layers of higher density, that is, in the above scenario, the layers arising from the disk rings.
@@
The interest of this approach is the alternative it offers to models in which the disk substructure is due to the presence of young planets. At the time of writing, no planetary-mass bodies have been detected in the HL Tau disk (Mullin24), nor in other structured disks, except in the systems PDS 70 (Keppler18) and WISPIT 2 (vanCapelleveen25), in both of which, however, the planets are located in large cavities 50–60 au wide. This point is therefore of crucial importance, as it indicates that magnetic effects must be taken into account in researches aimed at determining the timescale for the formation of the planets themselves.
@@
However, a number of issues need to be addressed before we can confirm the validity of the proposed association. First of all, the value of the magnetic lever arm assumed in this class of numerical models is generally 2. A similar simulation for , as suggested by our analysis, would be desirable. Secondly, the vertical size of the computational domain varies among the models, but in general the simulations do not reach the region at 103 au from the star in which we observe the wind structures with ALMA. An extension of numerical studies to large vertical scales would be needed, but it would come at the expense of a consistent computational effort. Finally, in these simulations, the disk structures are often seen to slowly move (Suriano19; Martel22), questioning the survival of their connection to the wind. However, the motion occurs on secular timescales, suggesting that the coherence of the magnetic flux tubes is probably not disrupted. Clearly, the applicability of the class of models described above to the observed cases needs to be tested with refined theoretical studies, which, however, are out of the scope of this work.
@@ 6 Conclusions
We investigated the properties of the molecular CO wind associated with the young HL Tau system by means of a detailed spectro-imaging analysis of the 12CO (2-1) line emission, observed with ALMA at 03 angular resolution in the context of the ALMA-DOT campaign (Podio20a; Garufi21). Our analysis confirms previous results derived from integrated intensity maps, but new insights on the structure of the flow come from the inspection of the individual projections of the position-position-velocity (PPV) datacube in the brighter and less perturbed redshifted lobe.
@@
A complex substructure is revealed in the velocity channel maps, constituted by concatenated bubble- and arc-shaped features that progressively increase in size and distance from the source with velocity. The conical shape of the wind appears to be the result of superposition of the limb-brightened borders of these features. The position-velocity (PV) diagram formed along the system axis presents a fan of separated almost linear traces opening from the source region, and with increasing velocity with offset from the star. The traces have strict spatial correspondence with the arcs in the channel maps. Finally, PV diagrams taken across the flow at stepped separations from the source show a series of concatenated ellipse-shaped features that correspond directly to the traces in the longitudinal PV diagram and in the channel maps. The ellipses are tilted with respect to the system axis, indicating rotation of the flow in the clockwise direction looking from the SW lobe toward the star, which matches the direction of the disk rotation.
@@
On the basis of these projections we infer a global structure of the emission in the PPV datacube consisting of a series of distinct concatenated paraboloid surfaces with the apex toward the source, with progressively larger aperture as the angle of their axis with respect to the velocity axis increases (cf. Fig. 7). In physical space, this translates into a finite number of coaxial, distinct, rotating nested gas shells. Ellipse fitting of the three directly identifiable structures in adjacent transverse PV diagrams allows a tomographic reconstruction of the morphology and kinematics of the outer three shells of the flow. We find smaller radii, higher poloidal and toroidal velocities, and steeper increases of the velocity with distance from the source for shells progressively closer to the system axis.
@@
This structure can be reproduced by a number of different wind acceleration mechanisms. In this paper, we focused on an analysis of the applicability of magnetohydrodynamic (MHD) disk winds (DW) accelerated from an extended region of the disk surface. We defer the comparison with other classes of models to future investigations.
@@
Under the DW hypothesis, we derive from the rotation properties of the flow that each examined shell is rooted in the disk within a few au of one of three adjacent disk rings, seen in the continuum emission of the dust, located at 58, 72, and 86 au from the star (ALMA15). However, because of the faintness of the examined flow substructures, we caution that the uncertainty on the footpoint radii is greater than the distance of the adjacent rings and gaps.
@@
We attempted a determination of the magnetic lever arm in the wind by comparing the observed trends of the wind parameters along the shells with the predictions of models of magnetized winds from either a laminar AD-dominated disk or a fully turbulent disk. The characteristics of the HL Tau wind turn out to be intermediate between the two types of solutions, with a value of . We discuss this finding in the context of the characteristics of the examined models. The retrieved launching region between 50 and 90 au would point to a wind from an AD-dominated outer disk, but solutions with such a high value of have yet to be found (current solutions reach = 2.76). A possibility would be to consider solutions for a higher level of AD than is usually assumed, which may be applicable if the wind itself screens the incoming ionizing cosmic rays by magnetic mirroring (Cleeves15). The high value of could also be justified by a magnetic field amplitude in the outer disk region larger than the value commonly adopted in these models as the ejection efficiency in the wind would be higher in this case.
@@
A magnetically driven disk wind with the above characteristics could sustain disk accretion at large scale by removing angular momentum in regions where the effective turbulence is suppressed because of weak internal ionization. In addition, the presence of distinct denser nested wind shells rooted in disk rings appears to support the results of recent non-ideal MHD numerical simulations for AD-dominated disks, which see the spontaneous formation of a ring–gap substructure in the disk connected with an inhomogeneous layered flow (e.g., Suriano19; Riols20), in a way alternative to the action of the yet elusive protoplanets. These aspects may have profound implications for the evolution of the disk and the process of planet formation.
@@
The inferred properties of the CO wind call for further studies of both observational and theoretical type. An analogous scrutiny of the fainter components and substructures in both the red- and blueshifted lobes in the same datacube, as well as the challenging estimate of the mass and angular momentum transported in each shell, requires the development of dedicated procedures, which is currently underway (Nony et al., in prep.). On the theoretical side, it would be interesting to consider the extension of the numerical simulations of inhomogeneous layered disk winds to the large scales probed by our observations. Then, a comparison of the retrieved results with other classes of wind models will be attempted, and will be the subject of future works. Finally, new observational studies for this and other systems, currently in progress, will allow us to analyze the dynamic relationship between the outer CO outflow, the coaxial H2 wind, and the atomic jet.
@@
These investigations are expected to provide further insight into the nature of the ejection phenomenon and finally enable us to quantify the feedback of the winds on the disk evolution.
@@ Acknowledgements.
This paper uses ALMA data 2018.1.01037.S (PI L. Podio). ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. We thank the anonymous referee for precious comments that led to improved results. The authors thank G. Lesur for fruitful discussions and for his support in using the self-similar MHD disk wind solutions that he has made available to the public at: https://github.com/glesur/PPDwind. The authors thank Z.-Y. Li, Y. Bai, T. Downes, F. Louvet, D. Fedele, and L. Testi for their helpful comments, and K. Rygl, V. Delabrosse, and U. Locatelli for their help with data processing and analysis. This work was supported by INAF-LG 2022 ”YSOs Outflows, Disks, and Accretion: towards a global framework for the evolution of planet-forming systems (YODA)” and in part by the NextGenerationEU funds within the National Recovery and Resilience Plan (NRRP), Mission 4 - Component 2 - From Research to Business (M4C2), Investment 3.1 - Project IR0000034 – ”STILES - Strengthening the Italian Leadership in ELT and SKA”. LP and CC acknowledge the projects PRIN-MUR 2020 - ”BEYOND-2p: Astrochemistry beyond the second period elements” (Prot. 2020AFB3FX); ASI-Astrobiologia 2023 - ”MIGLIORA-Modeling Chemical Complexity”, (F83C23000800005); INAF-GO 2024 - ”ICES: Tracking the history of ices from the cradles of planets to comets”; INAF-GO 2023 - ”PROTO-SKA-Exploiting ALMA data to study planet-forming disks: preparing the advent of SKA” (C13C23000770005); INAF-MG 2022 - ”Chemical Origins” (PI: L. Podio); financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title 2022JC2Y93 ”ChemicalOrigins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks” – CUP J53D23001600006 - Grant Assignment Decree No. 962 adopted on 30.06.2023 by the Italian Ministry of Ministry of University and Research (MUR).Appendix A Atlas of velocity channel maps
Figures 14 and 15 present an atlas of channel maps of CO (2-1) emission at VLSR from -3.0 km sto +28.6 km s, in steps of 0.4 km s. The panels at VLSR from +3.0 to +8.6 km sare centered on the source, to visualize the low-velocity emission in both lobes, while outside this range, emission is detected only in the blueshifted or in the redshifted lobe, so the maps cover the relevant quadrant, with the source positioned in one corner.
The emission in the blueshifted NE lobe is very faint. It consists of a large arc-shaped feature opening toward the NE seen at progressively larger distances from the source with increasing blueshifted velocities. A finger-like feature is identified at almost all velocities in about the same position north to the source and appears to be independent on the moving arc.
The redshifted SW lobe is distributed in two main components, a low-velocity wide-angle wind and a faster inner flow of apparent conical shape, with an opening angle decreasing with velocity. To quantify the variation of the aperture, we determined, for selected channels from 9 to 28 km sand separately for the SE and NW sides, the angle between the system axis and a line connecting the source and the position of the peak emission at the flow borders (identified by the contour profiles), at an offset of 45 and 75 from the source. The results of this analysis are shown in Figure 16. The half-opening angle decreases on average with velocity, from about 40 2 at 10 km sto 27 2 at 26 km s. The difference in value displayed in a given channel by the various curves reflects the deviations in the shape of the flow from a perfect cone.